OpenStudy (anonymous):

find the specified intercept for the graph of each equation 2x+9y=-9 x-intercept: ?

OpenStudy (anonymous):

Just make x=0 and you can compute the y-value for that. That is your x-intercept.

OpenStudy (anonymous):

so its zero

OpenStudy (anonymous):

No I am sorry, I was mistaken above. That will give you the y-intercept. To solve for the x-intercept you have to let y=0; \[2x+9y=-9\]Where y=0\[2x+9*0=-9\]and now you can divide both sides by 2\[2x=-9\]which would give you\[x = -\frac{ 9 }{ 2 } = -4.5\] And that is your x-intercept. For your y-intercept just repeat this procedure and make x=0.

OpenStudy (anonymous):

oh alright thanks